OpenStudy (anonymous):

help me plz. again w/ this one 12+2xlessthan or equal to 6 Thank you!

OpenStudy (anonymous):

The question is \(12+2x\le6\) and solve x...

OpenStudy (anonymous):

1. subtract 12 from both sides. 2. divide both sides by 2.

OpenStudy (anonymous):

12 + 12x <= 6 (subtract 12 from each side) 12 - 12 + 12x <= 6 - 12 12x <= 6 - 12 12x <= - 12 (divide by 12 to get x by itself) 12x/12 <= - 12/12 x <= - 1

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (anonymous):

np

OpenStudy (amistre64):

It is better to guide the user to an answer rather than doing all the work yourself. We discourage giving out answers without letting the asker at least take some part in the analysis - the aim of this site is to teach. Please refer to our code of conduct at http://openstudy.com/code-of-conduct for more information. Thank you.

OpenStudy (anonymous):

i got negative 3 ----------------- 12+2x>6 -12 -12 2x>-6 2x/2>-6/2 x>-3

OpenStudy (anonymous):

oops I wrote the problem wrong....srry

OpenStudy (anonymous):

it's okay!, I've got a clue from your answer! :D

OpenStudy (anonymous):

answer is - 3

OpenStudy (anonymous):

Yes . .. Thannk you so much, I also got x≤-3 :)